OpenStudy (mitu12):

solve the inequality -8+9x greater than or equal to 82.

OpenStudy (mitu12):

@KlOwNlOvE

OpenStudy (mitu12):

im stuck on this one part

OpenStudy (anonymous):

\[-8+9x \ge82\] is the same as \[9x-8\ge82\] what do you think we do first?

OpenStudy (mitu12):

add 8 form 82 and you get 100

OpenStudy (anonymous):

82+8=100?

OpenStudy (mathmale):

Add 8 TO 82 and get 90!

OpenStudy (mitu12):

\[9x \ge 100\]

OpenStudy (anonymous):

right operation wrong sum

OpenStudy (mitu12):

ohhh i added 18

OpenStudy (anonymous):

82+9=90

OpenStudy (anonymous):

oops 82+8=90*

OpenStudy (mitu12):

x greater than or equal to 10

OpenStudy (anonymous):

\[9x \ge90\]

OpenStudy (anonymous):

yes

OpenStudy (mitu12):

thank you all

OpenStudy (anonymous):

you're welcome